a path surrounds a plot of grass 45 ft. wide and 35ft. long. If the path is 6ft. wide, find the area [of the path.]
Answer by TimothyLamb(4379)

Ag = 45*35 = 1575 sq.ft.
Agp = (45+12)*(35+12) = (57)*(47) = 2679 sq.ft.
---
Ap = Agp - Ag = 2679 - 1575 = 1104 sq.ft.
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php